Respected Sir/Ma'am : I have a liitle problem with solving this problem and i am currently in Grade:9. The problem is as follows: If x=9-4*(5)^1/2 then find the value of : (x)^1/2-1/(x)^1/2 ;[(n)^1/2 which is Root:n)]. Thank You in advance Sir/Ma'am. Also i want to know if i have any doubts in the problem solving, how can i approach to you, imean how can i consult you back?


  1. Respected Sir/Ma'am :


       I have a liitle problem with solving this problem and i am currently in Grade:9. The problem is as follows:


If x=9-4*(5)^1/2 then find the value of :


(x)^1/2-1/(x)^1/2                       ;[(n)^1/2 which is Root:n)].


Thank You in advance Sir/Ma'am.


Also i want to know if i have any doubts in the problem solving, how can i approach to you, imean how can i consult you back?

Grade:11

2 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Neel,

Ans:- √x - 1/√x=(√(√x - 1/√x))²

                       =√(x+1/x- 2)

                             Putting the values and using 9-4√5 =1/(9+4√5) we get. The ans is=4

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.
win exciting gifts by answering the questions on Discussion Forum.

All the best Neel !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR

Vijay Luxmi Askiitiansexpert
13 years ago

Dear Neel,

√x - 1/√x=(√(√x - 1/√x))²

                       =√(x+1/x- 2)

                             Putting the values and using 9-4√5 =1/(9+4√5) we get. The ans is=4

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.
win exciting gifts by answering the questions on Discussion Forum.